Reports 1-1 of 1 Clear search Modify search
yohei.nishino - 22:39 Tuesday 17 May 2022 (20764) Print this report
Final check of OMC and OFI

Aso, Micahel, Nishino

We did a final check of OMC and OFI.


It seems there was a lot of junk light in the last experiment setup (20747). So we changed its procedure like this;

  1. We first maximized S-pol power with an HWP and polarizer and measured the transmissivity in the forward direction (T_f). (The polarizer was removed when measuring T_b)
  2. We set a polarizer maximizing its transmitted power and preserving its angle.
  3. Flip the OFI itself and the angle-preserved polarizer, and measure transmissivity In this setup, we cannot see the backward transmitted light by eyes, so we inserted an HWP in front of the QP (see figure), and slightly change the polarization angle to visualize the backward transmission.

Transmissivity = 1.662 [mW] / 1.708 [mW] = 97.3 % 

Extinction ratio = 10*log10(18.72 [mW] / 5.27 [μW]) = 35.5 dB


We applied first contact and cleaned the mirror surface. Then transmissivity and finesse are greatly improved from (20719).

Injection power = 22.3 mW

Transmissivity = 97.0 %

Reflectivity = 1.9 %

Loss = 1.1 %

Strage time = 1.27 μsec

Finesse = 813 

From these values, the loss for each mirror can be estimated as ~3.4 ppm.

Inspection in Mitaka is finished and we will pack and bring them to Kamioka next week.

Images attached to this report
Search Help